Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 18/7/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Hall Manager - water system change-over procedure (stands) - have laminated document ready to post (Colin/Josef), need more visible labelling of valves; stands
    • Check paper logbook for issues addressed there (SBoyd) - stands
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef)
    • think about the integration between LH2 and the rest of MICE : (Craig, Josef, Henry)
    • Implement weekly alarm review procedure for MOM (Steve): stands
    • Henry to show Colin beamstop issues
  • DC update
  • State of air conditioning (Colin)
  • Current alarm clearance plans (DC/Ajit)
  • Hydrogen system update and plans (JBoehm)
  • Cycle 2017 Issues
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
  • AOB

Minutes

  • Actions
    • Hall Manager - water system change-over procedure - have laminated document on MLCR wall - done, still need more visible labelling of valves
    • Check paper logbook for issues addressed there (SBoyd) - stands
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef)
    • think about the integration between LH2 and the rest of MICE : (Craig, Josef, Henry)
      • When LH2 present will lock off only the S Mezz. with access under DC control. Shifters should escalate issues via MOM/DC to experts.
    • Implement weekly alarm review procedure for MOM (Steve): stands
    • Henry to show Colin beamstop issues - superseded
  • DC update
    LH2 system cooldown and Neon test is in progress. PW down today as some issues with the sequences.
    LH2 Safety tour generated a list of minor electrical modifications; Richard Hale is not ATEX qualified. Need to find a person that can either do or supervise the work.
    Synch SP7: ISIS are dismantling/rebuilding the dipole adjacent to the Target - we must watch out for collateral damage
    We need to get Paul Masterson involved with the MICE water and A/C
    SSD insulating vacuum is gradually getting worse - needs investigation when Mark Tucker gets time
  • State of air conditioning (Colin)
    A/C: 2 is running but noisy; 3 is on but reports a "probe error"; 1 is off. Shouldn't run too many at the same time since this causes the compressors to start/stop leading to failure.
  • Current alarm clearance plans (DC/Ajit)
    Go through alarms; notify experts; get them fixed. Need to decide on a place to keep the list of (disabled) persistent alarms.
    AKurup is working on a script that identifies all presently-alarming PVs, and dumps a text file to be stored.
    Need now to treat all alarms as real; and follow up (check Archiver) to understand the issues
  • Hydrogen system update and plans (KLong)
    Need to work through all the control sequences with Neon to confirm they work.
    Had the first safety tour. Final tour sometime 14th August on, then can test with H2.
  • Cycle 2017 Issues
    The meeting went through the "Other" section of the list:
    • PHodgson to contact Paul Masterson re. how to contact A/C service company
    • Mismatched numbers of Triggers - need fresh data?
    • SBoyd to update shift checklist
    • A/C units 2 and 3 are now reading out the correct way round.
    • "U/S tracker" item: Henry complained that a significant number of issues had simply been cut out of the list; e.g. the U/S tracker problem (which had appeared between the expert-led and data runs) had not been spotted by the shifters since there were no reference plots for the Online Monitoring. Realistically, there will be no reference plots. But then that should be the close-out note in the list, so that people know what's going on (isn't the Online Monitoring irrelevant to shifters without them?). Presumably the Tracker group are aware of the failure mode and have taken steps to prevent a repeat.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB
    • Have asked to run @ 3V next cycle; needs a formal request. Should ask for Machine Physics time for beambump twiddling, etc.
    • Need to find the MICE van keys
    • DSA is now open and accessible (radiation survey OK). Dead GVA1 and old Manual Proton Absorber assembly in Hall to be moved to R9.
    • Ideally the PRY can now stay closed, but we will need a Geometry survey: arrange for September
    • Durga expected back in September

Updated by Nebrensky, Henry almost 5 years ago · 11 revisions